OpenStudy (anonymous):

Describe how you can use the graph of y=x2 to graph the given function. Help please y=x2-8

jimthompson5910 (jim_thompson5910):

To graph \(\Large y = x^2 - 8\), you simply shift the graph of \(\Large y = x^2\) 8 units down.

OpenStudy (anonymous):

can you share a drawing of what your talking about ?

jimthompson5910 (jim_thompson5910):

So here's y = x^2 

jimthompson5910 (jim_thompson5910):

now shift y = x^2 eight units down to get y = x^2 - 8